# Fast Matrix Multiplication via Ternary Meta Flip Graphs

### v1

@@ Abstract
Matrix multiplication optimization remains a fundamental challenge in computational mathematics. This work introduces a novel approach that discovers matrix multiplication schemes in the ternary field ($Z_{T}$), where coefficients are restricted to $\{-1,0,1\}$ to minimize naive additive complexity. The core of the method is a GPU-accelerated meta flip graph algorithm that maintains ternary safety through specialized arithmetic operations and sign symmetry breaking. Key results include new best ranks for the formats $4\times 5\times 12$, $5\times 6\times 10$, and $6\times 7\times 9$, the independent discovery of 32 schemes in $Z_{T}$ that match known optimal ranks (including 8 previously known only with rational coefficients), and 30 rank improvements in the binary field. The analysis of 164 known schemes shows that 92 can be implemented in $Z_{T}$, while 72 could not be found in the ternary field with current methods, defining the current boundaries of this approach. All software, results, and discovered schemes are provided as open-source.
Keywords Fast matrix multiplication $\cdot$ Flip graph $\cdot$ Ternary integer field $\cdot$ Tensor rank
@@ 1 Introduction
This work focuses on matrix multiplication in the ternary field ($Z_{T}$), using only the coefficients $\{-1,0,1\}$. This avoids costly multiplications by large numbers while keeping algorithms efficient. A new GPU-accelerated search algorithm was built to explore only ternary schemes, using custom arithmetic and sign symmetry breaking rules. The system also includes new operations – like project, extend, merge, product, and size swap – to efficiently explore the space of possible algorithms.
@@
  • Rediscovery of 32 schemes in the ternary field ($Z_{T}$) that match the best-known ranks, including 8 that were previously only known with rational coefficients ($Q$), and 30 rank improvements in the binary field ($Z_{2}$).
@@
  • The fields and rings of coefficients are denoted as:
    • $Z_{T}=\{-1,0,1\}$ – the ternary field;
    • $Z_{2}=\{0,1\}$ – the binary field;
@@
This notation distinguishes schemes using only simple coefficients ($Z_{T}$) from those requiring larger integers ($Z$) or fractions ($Q$).
@@ 2.1 Matrix Multiplication Complexity
While recent theoretical work has improved the asymptotic complexity to $O(n^{2.371339})$ (alman2025more), these methods have large constant factors that make them impractical for real use. For small matrices, researchers now focus on finding the exact minimum number of multiplications needed – called the rank – for specific matrix sizes like $3\times 3\times 3$ or $4\times 4\times 4$. These small algorithms are actually useful in real applications like deep learning and scientific computing.
@@ 2.2.1 SAT Solving
SAT solvers have been applied by encoding the problem as Boolean satisfiability instances. For small formats like $2\times 2\times 2$, SAT solvers can find solutions in seconds. However, for $3\times 3\times 3$ matrices, exhaustive SAT solving becomes infeasible, running for days without finding solutions.
Recent work has improved SAT-based approaches through symmetry breaking techniques, which help reduce the search space by eliminating redundant solutions (yang2024ruling). The practical solution also combines SAT with local search heuristics. By starting from randomly initialized coefficients, researchers have discovered more than 17,000 non-equivalent schemes for $3\times 3\times 3$ multiplication (heule2019local; heule2021new). This hybrid approach shows that while pure SAT solving faces scalability limits, guided search with SAT components remains valuable.
@@ 2.4 Additive Complexity
Besides multiplication count, the number of additions and subtractions (additive complexity) matters greatly in practice. Algorithms with minimal multiplications but many additions can be slower in real hardware (brent1970algorithms). This work focuses on finding algorithms that minimize both multiplications and naive additions by using only simple coefficients: -1, 0, and 1.
@@ 3.1.1 Example: a 2×2 by 2×3 Matrix Multiplication Scheme
3.2 The Ternary Field Approach
This work explores matrix multiplication schemes restricted to the ternary field $Z_{T}=\{-1,0,1\}$, which offer practical advantages for hardware implementation. While schemes with larger coefficients (2, 4, 9, etc.) or fractions may achieve theoretically optimal ranks, they introduce substantial additive complexity that reduces their practical utility. The ternary constraint enables efficient hardware representation using minimal resources, eliminating the need for expensive multiplication or division operations during coefficient application.
In the ternary system all coefficients in the tensors $U$, $V$ and $W$ are constrained to $Z_{T}$, ensuring the entire computation requires only additions and subtractions without scalar multiplications by coefficients other than $\pm 1$. This approach maintains the mathematical structure described in Section 3.1 while imposing the practical constraint that all coefficients must be in $\{-1,0,1\}$.
@@ 3.4.1 Overall Algorithm Structure
    Rank assessment: calculate current best ranks for every unique matrix format $(n_{1}\times n_{2}\times n_{3})$ across the population, establishing baseline performance metrics.
@@ 3.5 Ternary Safety Mechanisms
Implementing ternary-safe operations presents greater computational challenges than working in the binary field. In $Z_{2}$, matrix schemes can be efficiently represented using uint64_t bitmasks with arithmetic handled through fast XOR operations. However, the ternary field $Z_{T}=\{-1,0,1\}$ requires representing three distinct states per coefficient instead of two, making simple uint64_t representations impossible.
A simple way would be to store the ternary numbers in arrays of small integers. However, this method is much slower because it uses more memory and does fewer calculations at a time. To keep the fast performance of the binary method while also using ternary numbers, a new system was created that packs the numbers into a special compact format.
@@ 3.6 Sign symmetry breaking
The ternary field introduces significant complexity in flip operations due to the presence of both positive and negative coefficients. A critical challenge arises from the need to consider both exact equality and inverse equality when identifying valid flip candidates. For example, pairs like $u^{(i)}\otimes v^{(i)}\otimes w^{(i)}$ and $u^{(j)}\otimes v^{(j)}\otimes w^{(j)}$ where $u^{(i)}=-u^{(j)}$ are mathematically valid for flip operations but require additional handling.
@@ 3.7 Search Strategy Variants
This approach enables efficient discovery of ternary-constrained matrix multiplication schemes while maintaining the performance necessary for comprehensive exploration of the algorithm space.
@@ 3.8.1 Ternary Lifting Script
This tool converts binary field ($Z_{2}$) schemes to ternary field implementations by solving the sign assignment problem. Using OR-Tools (perron2025google) constraint programming, the tool assigns ±1 signs to non-zero coefficients while preserving Brent equation 1 validity. This approach is related to the constraint programming methods used to discover matrix multiplication schemes from scratch (deza2023fast). The implementation supports solving for multiple valid sign assignments.
@@ 4.2 New Best Ranks
The research established new best-known ranks for three matrix formats in the ternary field:
@@
4.3 Rediscoveries in the Ternary Field
A key result of this work is the independent discovery of 32 matrix multiplication schemes in the ternary field ($Z_{T}$) that match the best-known ranks from other fields. This includes 8 schemes that were previously only known with rational coefficients ($Q$). As shown in Table 1 and Table 2, these schemes cover multiple matrix formats and achieve mathematical correctness using only the hardware-efficient coefficients $\{-1,0,1\}$, proving that complex or fraction coefficients are not necessary for these algorithms.
@@ 4.3 Rediscoveries in the Ternary Field
Table 2: Other rediscovered schemes in the ternary field
@@ 4.5 Naive Additive Complexity Reduction
Beyond multiplicative complexity, the research focused on optimizing naive additive complexity, which significantly impacts practical performance in hardware implementations. Table 4 presents the results for 40 optimized schemes, showing substantial reductions in the number of addition operations required. The additive complexity metric used (non-zero coefficients - $2\timesrank$ - $m\times p$) provides a practical measure of implementation efficiency.
@@ 4.6 Ternary Schemes with Non-Optimal Ranks
The search also discovered several schemes in $Z_{T}$ that, while not achieving the absolute minimal known rank, represent progress toward ternary implementations. For formats where the optimal rank requires coefficients outside 1,0,1{-1,0,1}, these schemes provide ternary alternatives. As shown in Table 5, these results represent the current frontier for ternary-constrained matrix multiplication in these formats. The search for schemes that achieve both the minimal rank and ternary coefficients remains an active research direction.
@@ 5.1 Implications of Ternary Constraints
This work shows that matrix multiplication algorithms using only $\{-1,0,1\}$ coefficients can be just as good as those using larger numbers. The 32 schemes found in the ternary field and 3 new best ranks prove that complex coefficients are not needed for optimal performance. This challenges the common belief that large numbers are necessary for the best algorithms.
@@
The analysis of 164 known schemes reveals current limitations: while 58 schemes were already known in $Z_{T}$ and 32 more were rediscovered in this work, 72 schemes remain elusive in the ternary field. This indicates that ternary constraints are feasible for many but not all optimal algorithms, highlighting an important boundary for this approach.
@@ 6 Conclusion
The main achievement is proving that simple coefficients are enough to create highly efficient algorithms for many cases. The discovery of 3 new best ranks and 32 schemes in the ternary field that match known optimal ranks demonstrates this capability. However, analysis of 164 known schemes shows that while 92 can be implemented in $Z_{T}$ (57 previously known plus 35 newly discovered), 72 currently cannot, clearly defining the current boundaries of what’s possible with ternary coefficients.
@@ 7 Availability
    • New best-rank schemes in ternary field;
@@
    • Experimental results and performance data.

### v2

@@ Abstract
Matrix multiplication optimization remains a fundamental challenge in computational mathematics. This work introduces a novel approach that discovers matrix multiplication schemes whose coefficients are restricted to the set $\{-1,0,1\}$ (denoted $Z_{T}$), minimizing naive additive complexity for efficient hardware implementation. The core of the method is a GPU-accelerated meta flip graph algorithm that maintains ternary safety through specialized arithmetic operations and sign symmetry breaking. Key results include new best ranks for the formats $4\times 5\times 12$, $5\times 6\times 10$, and $6\times 7\times 9$, the independent discovery of 32 schemes in $Z_{T}$ that match known optimal ranks (including 8 previously known only with rational coefficients), and 30 rank improvements in the binary field. The analysis of 164 known schemes shows that 92 admit a ternary-coefficient implementation, while 72 could not be found under this constraint, defining the current boundaries of the approach. All software, results, and discovered schemes are provided as open-source.
Keywords Fast matrix multiplication $\cdot$ Flip graph $\cdot$ Ternary integer coefficient set $\cdot$ Tensor rank
@@ 1 Introduction
This work focuses on matrix multiplication schemes where all coefficients are restricted to the set $Z_{T}=\{-1,0,1\}$, referred to as ternary coefficients. This approach should not be confused with working over the finite field $\mathbb{F}_{3}$. Instead, it constrains coefficients for hardware efficiency while maintaining schemes over $\mathbb{Q}$ or $\mathbb{R}$. This avoids costly multiplications by large numbers while keeping algorithms efficient. A new GPU-accelerated search algorithm was built to explore only ternary schemes, using custom arithmetic and sign symmetry breaking rules. The system also includes new operations – like project, extend, merge, product, and size swap – to efficiently explore the space of possible algorithms.
@@
  • Rediscovery of 32 schemes using ternary coefficients ($Z_{T}=\{-1,0,1\}$) that match the best-known ranks, including 8 that were previously only known with rational coefficients ($Q$), and 30 rank improvements in the binary field ($Z_{2}$).
@@
  • The coefficient sets and rings are denoted as:
    • $Z_{T}=\{-1,0,1\}$ – the ternary coefficient set;
    • $Z_{2}={0,1}$ – the binary field;
@@
This notation distinguishes schemes using only simple ternary coefficients ($Z_{T}$) from those requiring arbitrary integers ($Z$) or rational numbers ($Q$).
@@ 2.1 Matrix Multiplication Complexity
While recent theoretical work has improved the asymptotic complexity to $O(n^{2.371339})$ (alman2025more), these methods have large constant factors that make them impractical for real use. For small matrices, researchers now focus on finding the exact minimum number of multiplications needed – called the rank – for specific matrix sizes like $3\times 3$ or $4\times 4$. These small algorithms are actually useful in real applications like deep learning and scientific computing.
@@ 2.2.1 SAT Solving
SAT solvers have been applied by encoding the problem as Boolean satisfiability instances. For small formats like $(2,2,2:7)$, SAT solvers can find solutions in seconds. However, for $(3,3,3:23)$ format, exhaustive SAT solving becomes infeasible, running for days without finding solutions.
Recent work has improved SAT-based approaches through symmetry breaking techniques, which help reduce the search space by eliminating redundant solutions (yang2024ruling). The practical solution also combines SAT with local search heuristics. By starting from randomly initialized coefficients, researchers have discovered more than 17,000 non-equivalent $(3,3,3:23)$ schemes (heule2019local; heule2021new). This hybrid approach shows that while pure SAT solving faces scalability limits, guided search with SAT components remains valuable.
@@ 2.4 Additive Complexity
Besides multiplication count, the number of additions and subtractions (additive complexity) matters greatly in practice. Algorithms with minimal multiplications but many additions can be slower in real hardware (brent1970algorithms). This work focuses on finding algorithms that minimize both multiplications and naive additions by using only simple coefficients: 1-1, 0, and $1$. The approach maximizes the number of zero coefficients in the scheme tensors, directly reducing the naive addition count. This provides a foundation for further optimization through common subexpression elimination (maartensson2025number), which can significantly reduce the actual number of operations in implementation. A detailed treatment of advanced addition minimization techniques will be presented in a separate study to maintain focus on the core rank optimization problem.
@@ 3.1.1 Example: a 2×2 by 2×3 Matrix Multiplication Scheme
3.2 The Ternary Coefficients Approach
This work explores matrix multiplication schemes restricted to the ternary coefficient set $Z_{T}=\{-1,0,1\}$, which offer practical advantages for hardware implementation. While schemes with larger coefficients (2, 4, 9, etc.) or fractions may achieve theoretically optimal ranks, they introduce substantial additive complexity that reduces their practical utility. The ternary constraint enables efficient hardware representation using minimal resources, eliminating the need for expensive multiplication or division operations during coefficient application.
In this system, all coefficients in the tensors $U$, $V$, and $W$ are constrained to $Z_{T}$, ensuring the entire computation requires only additions and subtractions without scalar multiplications by coefficients other than $\pm 1$. This approach maintains the mathematical structure described in Section 3.1 while imposing the practical constraint that all coefficients must be in $\{-1,0,1\}$.
@@ 3.4.1 Overall Algorithm Structure
    Rank assessment: calculate current best ranks for every unique matrix format $(m\times n\times p)$ across the population, establishing baseline performance metrics.
@@ 3.5 Ternary Safety Mechanisms
Implementing operations that preserve the ternary coefficient constraint presents greater computational challenges than working in the binary field. In $Z_{2}$, matrix schemes can be efficiently represented using uint64_t bitmasks with arithmetic handled through fast XOR operations. However, the ternary set $Z_{T}=\{-1,0,1\}$ requires representing three distinct states per coefficient instead of two, making simple uint64_t representations impossible.
A straightforward approach would store ternary coefficients in arrays of small integers. However, this method incurs substantial performance penalties due to increased memory bandwidth requirements and reduced computational density. To maintain the high performance characteristics of binary field implementations while supporting ternary constraints, a specialized packed representation was developed that efficiently encodes the three coefficient states.
@@ 3.6 Sign symmetry breaking
The ternary coefficient set introduces significant complexity in flip operations due to the presence of both positive and negative coefficients. A critical challenge arises from the need to consider both exact equality and inverse equality when identifying valid flip candidates. For example, pairs like $u^{(i)}\otimes v^{(i)}\otimes w^{(i)}$ and $u^{(j)}\otimes v^{(j)}\otimes w^{(j)}$ where $u^{(i)}=-u^{(j)}$ are mathematically valid for flip operations but require additional handling.
@@
3.7 Isotropy Invariants and Novelty Verification
To characterize discovered schemes and verify they represent new tensor decompositions rather than points in known orbits under the de Groote isotropy group, three types of invariants are computed. For each scheme $(U,V,W)$ of rank $r$, the following invariants are calculated:
  • Type invariant: following the FMM Catalogue (sedoglavic2023yet), the polynomial
    $\sum_{i=1}^{r}X^{\text{rank}U^{(i)}}Y^{\text{rank}V^{(i)}}Z^{\text{rank}W^{(i)}}$
    is computed, where $\text{rank}(M)$ denotes the classical matrix rank of $M$.
  • Symmetric polynomial invariant: following the "3x3 Matrix Multiplication Solution Repository" (heule2019local), the symmetrized polynomial
    $f(x,y,z)=\sum_{\pi\in S_{3}}\pi\left(\sum_{i=1}^{r}x^{\text{rank}U^{(i)}}y^{\text{rank}V^{(i)}}z^{\text{rank}W^{(i)}}\right)$
    is computed, where the permutations $\pi\in S_{3}$ act on the variables $x,y,z$.
  • Rank sum polynomial: additionally, the polynomial
    $g(\omega)=\omega^{\sum_{i=1}^{r}\text{rank}U^{(i)}}+\omega^{\sum_{i=1}^{r}\text{rank}V^{(i)}}+\omega^{\sum_{i=1}^{r}\text{rank}W^{(i)}}$
    is computed.
These invariants are computed for all newly discovered schemes and compared against those of known decompositions. The complete invariant data, which is too extensive to include in the main text, is provided within the accompanying scheme files in the open-source repository. This analysis confirms that the ternary-constrained schemes discovered via the flip graph approach represent genuinely new points in the space of matrix multiplication algorithms.
3.8 Initial Scheme Sources
The search process was initialized using matrix multiplication schemes from established sources:
  • FMM Catalogue (sedoglavic2023yet);
  • AlphaTensor (fawzi2022discovering);
  • AlphaEvolve (novikov2025alphaevolve);
  • Flip Graph (kauers2023flip);
  • Adaptive Flip Graph (arai2024adaptive);
  • Symmetric Flip Graph (moosbauer2025flip);
  • Meta Flip Graph (kauers2025exploring).
From these sources, only schemes already employing ternary coefficients (1,0,1{-1,0,1}) were selected for initialization. This approach ensured the search process began within the feasible space of ternary-constrained implementations while building upon existing research.
@@ 3.9 Search Strategy Variants
This methodology enables efficient discovery of matrix multiplication schemes with ternary coefficients while maintaining the performance required for comprehensive exploration of the algorithm space.
@@ 3.10.1 Ternary Lifting Script
This tool converts binary field ($Z_{2}$) schemes to implementations with ternary coefficients by solving the sign assignment problem. Using OR-Tools (perron2025google) constraint programming, the tool assigns ±1 signs to non-zero coefficients while preserving Brent equation 1 validity. This approach is related to the constraint programming methods used to discover matrix multiplication schemes from scratch (deza2023fast). The implementation supports solving for multiple valid sign assignments.
@@ 4.2 New Best Ranks
The research established new best-known ranks for three matrix formats using ternary coefficients:
@@
4.3 Rediscoveries with Ternary Coefficients
A key result of this work is the independent discovery of 32 matrix multiplication schemes using ternary coefficients ($Z_{T}$) that match the best-known ranks from other fields. This includes 8 schemes that were previously only known with rational coefficients ($Q$). As shown in Table 1 and Table 2, these schemes cover multiple matrix formats and achieve mathematical correctness using only the hardware-efficient coefficients $\{-1,0,1\}$, proving that complex or fraction coefficients are not necessary for these algorithms.
@@ 4.3 Rediscoveries with Ternary Coefficients
Table 2: Other rediscovered schemes with ternary coefficients
@@ 4.5 Naive Additive Complexity Reduction
Beyond multiplicative complexity, the research focused on optimizing naive additive complexity, which significantly impacts practical performance in hardware implementations. This optimization, achieved by maximizing zero coefficients in the scheme tensors, provides a foundation for further reduction through common subexpression elimination. Table 4 presents the results for 40 optimized schemes, showing substantial reductions in the number of addition operations required. The additive complexity metric used (non-zero coefficients - $2\times r-m\times p$) provides a practical measure of implementation efficiency.
@@ 4.6 Ternary Schemes with Non-Optimal Ranks
The search also discovered several schemes using ternary coefficients that, while not achieving the absolute minimal known rank, represent progress toward efficient implementations. For formats where the optimal rank requires coefficients outside $\{-1,0,1\}$, these schemes provide ternary alternatives. As shown in Table 5, these results represent the current frontier for matrix multiplication with ternary coefficients in these formats. The search for schemes that achieve both the minimal rank and ternary coefficients remains an active research direction.
@@ 5.1 Implications of Ternary Constraints
This work shows that matrix multiplication algorithms using only $\{-1,0,1\}$ coefficients can be just as good as those using larger numbers. The discovery of 32 schemes with ternary coefficients matching known optimal ranks, plus 3 new best ranks, proves that complex coefficients are not necessary for optimal performance in many cases. This finding challenges the conventional assumption that large integer or rational coefficients are essential for achieving the best algorithmic performance.
@@
The analysis of 164 known schemes reveals current limitations: while 58 schemes were already known with ternary coefficients and 32 more were rediscovered in this work, 72 schemes remain elusive under the ternary constraint. This indicates that ternary coefficients are feasible for many but not all optimal algorithms, highlighting an important boundary for this approach.
@@ 6 Conclusion
The main achievement demonstrates that simple coefficients are sufficient to create highly efficient algorithms for many cases. The discovery of 3 new best ranks under ternary constraints, combined with 32 schemes that match previously known optimal ranks using only ternary coefficients, confirms this capability. However, analysis of 164 known schemes shows that while 92 can be implemented with ternary coefficients (57 previously known plus 35 newly discovered), 72 currently cannot be found under this constraint, clearly defining the current boundaries of what’s possible with ternary coefficients.
@@ 7 Availability
    • New best-rank schemes using ternary coefficients;
@@
    • Complete experimental results and data.